OpenStudy (anonymous):

Suppose that you were saving money over 5 years to use in a purchase later. You have $1000 to put in the savings. After surveying several banks for savings plans, you found these options. "A" stands for the amount you will have in the bank after x years. Option A: Your money would receive simple interest at the end of 5 years. The formula is A = 1000 + 1000(0.05) x. Option B: Your money will be compounded continuously. The formula is A = (1000)(2.71828)(0.05 x). Option C: You will invest in a CD (Certificate of Deposit) compounded yearly.

OpenStudy (anonymous):

The formula is A = (1000)(1.05) x. Option D: Buy a US Savings Bond. Buy at $1000 now. When cashed in 5 years later, the amount will be $1267. Calculate each of these. You can find this by substituting x = 5 in each formula. Note that there is no formula in the US Savings Bonds option, just the amount.

OpenStudy (anonymous):

simple interest? compound interest? continually interest?

OpenStudy (anonymous):

1. Show the calculations for Options A, B, and C. 2. What is the order of Options A, B, C, and D, listing the option which gives the greatest amount at the end of 5 years to the least? 3. Which option gave you the greatest amount at the end of the five years? 4. What was the amount you calculated with this option?

OpenStudy (anonymous):

Interest means more money; therefore, continually interest, option B would be best

OpenStudy (anonymous):

i'm so confused with this entire thing .. those 4 things above ^ are what i have to do and i don't get it .

OpenStudy (phi):

type this into the google search window for A: 1000 + 1000*(0.05)*5=

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

i got 1250

OpenStudy (phi):

not sure about your equation for B. But for C (1000)(1.05)^5 I assume x is the exponent?

OpenStudy (anonymous):

i think so

OpenStudy (phi):

type that expression in google to evaluate it.

OpenStudy (phi):

I think option B's expression is (1000)*(2.71828)^(0.05*5)= type that into google.

OpenStudy (phi):

the ^ means power. for example 3^4 means 3*3*3*3 (three times itself 4 times)

OpenStudy (anonymous):

oh alright(: thank you ! i think i can try and get it form here now (: thanks SOOOOO much for your help (: